# revtex4-1Repair the float

### v4

@@ Abstract
How does quantum chaos lead to rapid scrambling of information as well as errors across a system when one introduces perturbations in the dynamics? What are its consequences for the reliability of quantum simulations and quantum information processing? We employ continuous measurement quantum tomography as a paradigm to study these questions. The measurement record is generated as a sequence of expectation values of a Hermitian observable evolving under repeated application of the Floquet map of the quantum kicked top. Interestingly, we find that the reconstruction fidelity initially increases regardless of the degree of chaos or the strength of perturbations in the dynamics. For random states, when the measurement record is obtained from a random initial observable, the subsequent drop in the fidelity obtained is inversely correlated to the degree of chaos in the dynamics. More importantly, this also gives us an operational interpretation of Loschmidt echo for operators by connecting it to the performance of quantum tomography. We define a quantity to capture the scrambling of errors, an out-of-time-ordered correlator (OTOC) between two operators under perturbed and unperturbed system dynamics that serves as a signature of chaos and quantifies the spread of errors. Our results demonstrate not only a fundamental link between Loschmidt echo and scrambling of errors, as captured by OTOCs, but that such a link can have operational consequences in quantum information processing.
@@
The central goal of quantum chaos is to inform us about the properties of quantum systems whose classical counterpart is chaotic. How does chaos manifest in the quantum world, and what notions of complexity might be suitable to quantify it? Quantum theory comes with another layer of complexity hitherto unknown in our classical description of reality, the Hilbert space, which is a big space caves1996quantum . Quantum theory permits the state of the system to be any vector in this space, even permitting a coherent superposition of possibilities considered mutually exclusive in the classical world. Therefore, while classically chaotic dynamics generate classical information in the form of classical trajectories, quantum chaotic dynamics generate quantum information in the form of pseudo-random vectors in the Hilbert space. These states typically have a high entropy which is calculated according to a fixed fiducial basis wootters1990random ; bengtsson2017geometry . Vigorous thrust in the understanding of quantum many-body dynamical systems through dynamically generated entanglement  miller1999signatures ; bandyopadhyay2002testing ; wang2004entanglement ; trail2008entanglement ; furuya1998quantum ; lakshminarayan2001entangling ; seshadri2018tripartite and quantum correlations madhok2015signatures ; madhok2018quantum , deeper studies in the ergodic hierarchy of quantum dynamical systems gomez2014towards ; bertini2019exact ; aravinda2021dual ; vikram2022dynamical have been some recent milestones. Also, the out-of-time-ordered correlators (OTOCs) that attempt to capture operator growth and scrambling of quantum information have been very useful as a probe for chaos in quantum systems maldacena2016bound ; swingle2016measuring ; hashimoto2017out ; kukuljan2017weak ; swingle2018unscrambling ; wang2021quantum ; sreeram2021out ; varikuti2022out . These, coupled with the traditional approach to studies of level statistics haake1991quantum and Loschmidt echo peres1984stability ; peres1997quantum ; Goussev2012loschmidt ; gorin2006dynamics and complemented by the ability to coherently control and manipulate many-body quantum systems in the laboratory gong2001coherent ; gong2005quantum ; brif2010control ; smith2013quantum ; mirkin2021quantum , have brought us to a fork in our path. On the one hand, this is a harbinger of the possibility of building quantum simulators, an important milestone in our quest for the holy grail - a many-body quantum computer. On the other hand, the same properties that make quantum systems generate complexity will make them sensitive to errors that naturally occur in implementing many-body Hamiltonians.
@@
Quantum tomography gives us a window to study sensitivity to errors in quantum simulations of chaotic Hamiltonians lloyd1996universal ; johnson2014what . Quantum tomography uses the statistics of measurement records on an ensemble of identical systems in order to make the best estimate of the actual state $\rho_{0}$. We consider continuous weak measurement tomography protocol silberfarb2005quantum ; smith2006efficient ; riofrio2011quantum ; smith2013quantum , and the time series of operators can be generated by the Floquet map of a quantum dynamical system to investigate the role of chaos on the information gain in tomography madhok2014information ; sreeram2021quantum ; sahu2022effect . Here we unify two quantifiers of quantum chaos, namely Loschmidt echo and OTOC, through scrambling of errors in continuous weak measurement tomography as illustrated in Fig. 1.
An ensemble of $N$ identical systems $\rho^{\otimes N}_{0}$ undergo separable time evolution by a unitary $U(t)$. A weakly coupled probe will generate the measurement record by performing weak continuous measurement of an observable $\mathcal{O}$. For sufficiently weak coupling, the randomness of the measurement outcomes is dominated by the quantum noise in the probe rather than the measurement uncertainty, i.e., the projection noise. In this case, the quantum backaction is negligible, and the state remains approximately separable. Thus, we get the stochastic measurement record
@@
where ${\bf C=(\tilde{\mathcal{O}}^{T}\tilde{\mathcal{O}}})^{-1}$ is the covariance matrix and the inverse is Moore-Penrose pseudo inverse ben2003generalized in general. The estimated Bloch vector ${\bf r}_{ML}$ may not always represent a physical density matrix with non-negative eigenvalues because of the finite signal-to-noise ratio. Therefore we impose the constraint of positive semidefiniteness on the reconstructed density matrix and obtain the physical state closest to the maximum-likelihood estimate. To do this, we employ a convex optimization procedure vandenberghe1996semidefinite where the final estimate of the Bloch vector $\bf\bar{r}$ is obtained by minimizing the argument
@@
The above description represents an ideal scenario where the experimentalist has complete knowledge of the true dynamics (which is symbolized as unprimed variables describing the observables, $\mathcal{O}_{n}$, and covariance matrix, $\bf C$, thus generated over time) and they can properly reconstruct the state using Eq. (4). However, in reality, one never knows the true underlying dynamics, and there is always a departure from the ideal case due to inevitable errors and perturbations to the true dynamics. Thus, the experimentalist, oblivious to these, models their estimation using a covariance matrix, ${\bf C^{\prime}=(\tilde{\mathcal{O^{\prime}}}^{T}\tilde{\mathcal{O^{\prime}}}})^{-1}$. Here the primed variables represent the experimentalist’s knowledge of the dynamics in the laboratory, and as a result, they end up reconstructing an incorrect state as $\bar{\rho}^{\prime}$ from
@@
Figure 2: Effect of perturbation on tomography quantified by different metrics as a function of time with an increase in the level of chaos. The kicked top Floquet map of true (perturbed) dynamics $U_{\tau}=e^{-i(\lambda+\delta\lambda)J^{2}_{z}/2J}e^{-i\alpha J_{x}}$ with $\delta\lambda=0.01$, and ideal (unperturbed) dynamics $U^{\prime}_{\tau}=e^{-i\lambda J^{2}_{z}/2J}e^{-i\alpha J_{x}}$ generate the time series of operators for a spin $j=10$ and fixed $\alpha=1.4$. (a) Average reconstruction fidelity $\langle\mathcal{F}\rangle$ of the state $\bar{\rho}^{\prime}$ derived from Eq. (6) relative to the actual state $\left|\psi_{0}\right\rangle$, where the average is taken over 100 Haar random states. (b) The operator Loschmidt echo $\mathcal{F}_{\mathcal{O}}$ between two operators. (c) The quantum relative entropy $\mathcal{D}_{KL}$ of regularized operator evolved under unperturbed dynamics to the operator evolved under perturbed dynamics. (d) The operator incompatibility $\mathcal{I}_{\mathcal{O}}$ quantifies the scrambling of errors.
@@
The connection between chaos and information gain depends on the localization properties of the state, i.e. their inverse participation ratio, the degree of chaos, as well as how well the state is aligned with time-evolved measurement observables sahu2022effect . Therefore, to study the effect of the degree of chaos on the performance of noisy tomography purely, we consider random initial states measured via random initial observables (generated by rotating $J_{x}$ through random unitary) picked from the appropriate Haar measure. We apply our reconstruction protocol on an ensemble of 100 random pure states sampled from the Haar measure on SU$(d)$, where $d=2j+1=21$. We choose one random initial observable and generate the measurement record from the repeated application of the Floquet map of the quantum kicked top. The fidelity of the reconstructed state $\bar{\rho}^{\prime}$ obtained from Eq. (6) is determined relative to the actual state $\left|\psi_{0}\right\rangle$, $\mathcal{F}=\left\langle\psi_{0}\right|\bar{\rho}^{\prime}\left|\psi_{0}\right\rangle$ as a function of time. We notice that the reconstruction fidelity increases in the beginning despite the errors, and after a certain period of time, it starts decaying. The rise in fidelity during the initial time period is because any information, even if partially inaccurate, about a completely unknown random state offsets the presence of errors in its estimation. However, as time progresses, the effect of errors becomes significant. Beyond a certain time, we observe a decline in fidelity as the dynamics continues to accumulate errors that dominate the archive of information present in the measurement record. Most interestingly, the rate of this fidelity decay is inversely correlated with the degree of chaos in the dynamics. This is the analog of an interplay between the rapid information gain due to Lyapunov divergence, a “quantum” analog of the classical KS entropy, and Loschmidt echo leading to errors that cause fidelity decay.
@@
The operator Loschmidt echo that captures the overlap of the operators $\mathcal{O}_{n}$ and $\mathcal{O}^{\prime}_{n}$, decays with time. We can see from Fig. 2b that the operator Loschmidt echo decays much slower when the dynamics is chaotic than when it is regular. This behavior of the operator Loschmidt echo correlates positively with the rate of drop in reconstruction fidelity as demonstrated in Fig. 2a and Fig. 2b. The greater the distance between the operators at a given time, the greater the difference between the expectation values with respect to the state and the archive of the measurement record obtained through the time series. Our results give an operational interpretation of the operator Loschmidt echo by connecting it to a concrete physical task of continuous measurement quantum tomography. This also points to a beneficial way to probe these quantities in experiments using current techniques. Quantum relative entropy is a measure of distance between two quantum states. Here we use this metric to measure the distance between two operators $\mathcal{O}_{n}$ and $\mathcal{O}^{\prime}_{n}$. To treat both observables as density operators, we regularize them as follows. We construct a positive operator from an observable by retaining its eigenvectors and taking the absolute value of its eigenvalues. To normalize this operator, we divide it by its trace [see Methods for details]. Now we can determine the quantum relative entropy
$\mathcal{D}_{KL}(\rho_{\mathcal{O}_{n}}||\rho_{\mathcal{O}^{\prime}_{n}})={\rm{Tr}}\,({\rho_{\mathcal{O}_{n}}}(\log\rho_{\mathcal{O}_{n}}-\log\rho_{\mathcal{O}^{\prime}_{n}})),$ (8)
@@
This form of error scrambling is very general for dynamics with time-dependent as well as time-independent Hamiltonian [see Methods for details]. The growth of OTOC has been studied extensively as a quantifier for information scrambling under chaotic dynamics maldacena2016bound ; swingle2016measuring ; hashimoto2017out ; kukuljan2017weak ; swingle2018unscrambling ; wang2021quantum ; sreeram2021out ; varikuti2022out . Similarly, growth of $\mathcal{I}_{\mathcal{O}}$ implies scrambling of errors with time. It is apparent from Fig. 2d that the rate of error scrambling decreases with an increase in the value of the chaoticity parameter $\lambda$. This signifies that the measurement record is less affected by the error in the dynamics when one approaches a greater extent of chaos. In Eq. (6), the measurement record $\bf M$ is obtained from the true (perturbed) dynamics, but the covariance matrix $\bf C^{\prime}$, and $\tilde{\mathcal{O}^{\prime}}$ are determined from the experimentalist’s version of the dynamics (ideal or unperturbed). Thus, a higher rate of error scrambling for regular dynamics leads to a faster decay of reconstruction fidelity as the measurement record is more vulnerable. How errors scramble across a chaotic system, as given by Eq. (9), is itself an interesting quantifier of quantum chaos. Here we notice the correlation between scrambling of errors as captured by the incompatibility between the operator and its time evolution through the error unitary in Eq. (9) and operator Loschmidt echo, as viewed from the lens of quantum tomography under chaotic dynamics. This links two fundamental quantifiers of quantum chaos, complements findings in zurek2020information and provides a different but more intuitive connection.
We find dynamical signatures of chaos that quantify the scrambling of errors across a many-body quantum system that has consequences on the performance of quantum information and simulation protocols. We also give an operational interpretation of the operator Loschmidt echo by connecting it to the growth of distance between operators evolved in continuous measurement quantum tomography. Our results linking Loschmidt echo, error scrambling, and OTOCs will be helpful to the condensed matter community as well and in addressing broader issues involving non-integrable quantum systems pandey2020adiabatic .
Our work paves the way for further studies in the performance of quantum simulations under inadvertent noise. In the era of noisy, intermediate-scale quantum (NISQ) devices preskill2018quantum , the accuracy of an analog quantum simulator will decay after just a few time steps. The reliability of such analog quantum simulators is highly questionable even for state-of-the-art architecture when it is likely to exhibit quantum chaos hauke2012can ; lysne2020small . On the contrary, the digital quantum simulation is often associated with the inherent Trotter errors heyl2019quantum because of the discretization of the time evolution of a quantum many-body system as a sequence of quantum gates. Thus, a better understanding of errors in simulating many-body quantum systems and information processing protocols that exploit such rich dynamics is paramount. These signatures of chaos can be further explored using state-of-the-art experimental techniques involving cold atoms interacting with lasers and magnetic fields chaudhury2009quantum . In future work, we hope to further build upon our results to develop quantum analogs of the “classical shadowing lemma” that guarantee a true classical trajectory in the neighbourhood of any arbitrary simulated trajectory of a chaotic system in the presence of truncation errors due to finite precision anosov1967geodesic ; grebogi1990shadowing ; sauer1991rigorous ; sauer1997long ; vanivcek2004dephasing .
We are grateful to Arul Lakshminarayan for useful discussions. We thank Sreeram PG for helpful discussions. We acknowledge BS Datta Vikas for his inputs during the initial stage of this work. The authors would like to thank HPCE, IIT Madras, for providing the computational facility for numerical simulations. This work was supported in part by grant DST/ICPS/QusT/Theme-3/2019/Q69 and New faculty Seed Grant from IIT Madras. The authors were supported, in part, by a grant from Mphasis to the Centre for Quantum Information, Communication, and Computing (CQuICC) at IIT Madras.
Supplementary Material
Relative entropy calculation
While calculating the relative entropy, we regularize the operator in the following way. First, we diagonalize the time-evolved operator $\mathcal{O}_{n}$ as
$\mathcal{O}_{n}=V_{n}D_{n}{V_{n}}^{\dagger},$ (10)
where $V_{n}$ is the unitary matrix that diagonalizes $\mathcal{O}_{n}$. In the second step, we take the absolute value of the eigenvalues of $D_{n}$ and divide it by its trace to get $\tilde{D}_{n}=|D_{n}|/{\rm{Tr}}\,(|D_{n}|)$. We, then, construct a positive operator with unit trace that behaves as a density matrix while keeping the eigenvectors of the observable $\mathcal{O}_{n}$ intact as
$\rho_{\mathcal{O}_{n}}=V_{n}\tilde{D}_{n}{V_{n}}^{\dagger}.$ (11)
Derivation of Error scrambling expression
@@
The operator incompatibility we have computed for the operators evolved through unperturbed and perturbed dynamics. This form is very general, and we do not assume any specific form of the error or any particular form of Hamiltonian (whether time-dependent or time-independent), as we can see below. Let the operators at time $t$ be $\mathcal{O}(t)=U^{\dagger}(t)\mathcal{O}U(t)$ and $\mathcal{O^{\prime}}(t)=U^{\prime\dagger}(t)\mathcal{O}U^{\prime}(t)$, where $U(t)$ is the unitary for the unperturbed dynamics and $U^{\prime}(t)$ is the unitary for perturbed dynamics. Thus, the operator incompatibility is
@@
Now one can simplify this expression using the Hermiticity property of the physical observables $\mathcal{O}(t)=\mathcal{O}^{\dagger}(t)$.
@@
We can further simplify the above expression by exploiting the cyclic property of trace operation and the property of a unitary operator $U^{\dagger}U=UU^{\dagger}=\mathbb{1}$. Let us consider the first term in the above expression

### v5

@@ Abstract
How does quantum chaos lead to rapid scrambling of information as well as errors across a system when one introduces perturbations in the dynamics? What are its consequences for the reliability of quantum simulations and quantum information processing? We employ continuous measurement quantum tomography as a paradigm to study these questions. The measurement record is generated as a sequence of expectation values of a Hermitian observable evolving under repeated application of the Floquet map of the quantum kicked top. Interestingly, we find that the reconstruction fidelity initially increases regardless of the degree of chaos or the strength of perturbations in the dynamics. For random states, when the measurement record is obtained from a random initial observable, the subsequent drop in the fidelity obtained is inversely correlated to the degree of chaos in the dynamics. More importantly, this also gives us an operational interpretation of Loschmidt echo for operators by connecting it to the performance of quantum tomography. We define a quantity to capture the scrambling of errors, an out-of-time-ordered correlator (OTOC) between two operators under perturbed and unperturbed system dynamics that serves as a signature of chaos and quantifies the spread of errors. Our results demonstrate not only a fundamental link between Loschmidt echo and scrambling of errors, as captured by ”error OTOCs”, but that such a link can have operational consequences in quantum information processing.
I Introduction
@@
The central goal of quantum chaos is to inform us about the properties of quantum systems whose classical counterpart is chaotic. How does chaos manifest in the quantum world, and what notions of complexity might be suitable to quantify it? Quantum theory comes with another layer of complexity hitherto unknown in our classical description of reality, the Hilbert space, which is a big space caves1996quantum . Quantum theory permits the state of the system to be any vector in this space, even permitting a coherent superposition of possibilities considered mutually exclusive in the classical world. Therefore, while classically chaotic dynamics generates classical information in the form of classical trajectories, quantum chaotic dynamics generate quantum information in the form of pseudo-random vectors in the Hilbert space. These states typically have a high entropy which is calculated according to a fixed fiducial basis wootters1990random ; bengtsson2017geometry . Vigorous thrust in the understanding of quantum many-body dynamical systems through dynamically generated entanglement  miller1999signatures ; bandyopadhyay2002testing ; wang2004entanglement ; trail2008entanglement ; furuya1998quantum ; lakshminarayan2001entangling ; seshadri2018tripartite and quantum correlations madhok2015signatures ; madhok2018quantum , deeper studies in the ergodic hierarchy of quantum dynamical systems gomez2014towards ; bertini2019exact ; aravinda2021dual ; vikram2022dynamical have been some recent milestones. Also, the out-of-time-ordered correlators (OTOCs) that attempt to capture operator growth and scrambling of quantum information have been very useful as a probe for chaos in quantum systems maldacena2016bound ; swingle2016measuring ; hashimoto2017out ; kukuljan2017weak ; swingle2018unscrambling ; wang2021quantum ; sreeram2021out ; varikuti2022out . These, coupled with the traditional approach to studies of level statistics haake1991quantum and Loschmidt echo peres1984stability ; peres1997quantum ; Goussev2012loschmidt ; gorin2006dynamics and complemented by the ability to coherently control and manipulate many-body quantum systems in the laboratory gong2001coherent ; gong2005quantum ; brif2010control ; smith2013quantum ; mirkin2021quantum , have brought us to a fork in our path. On the one hand, this is a harbinger of the possibility of building quantum simulators, an important milestone in our quest for the holy grail- a many-body quantum computer. On the other hand, the same properties that make quantum systems generate complexity will make them sensitive to errors that naturally occur in implementing many-body Hamiltonians.
@@
Quantum tomography gives us a window to study sensitivity to errors in quantum simulations of chaotic Hamiltonians lloyd1996universal ; johnson2014what . Quantum tomography uses the statistics of measurement records on an ensemble of identical systems in order to make the best estimate of the actual state $\rho_{0}$. We consider continuous weak measurement tomography protocol silberfarb2005quantum ; smith2006efficient ; riofrio2011quantum ; smith2013quantum ; PhysRevLett.95.030402 ; Riofrío_2011 ; PhysRevA.90.032113 ; PhysRevLett.121.130404 ; PhysRevLett.97.180403 ; PhysRevA.87.030102 ; PhysRevLett.111.160401 ; PhysRevLett.114.090403 ; hacohen2016quantum ; PhysRevLett.115.180407 , and the time series of operators can be generated by the Floquet map of a quantum dynamical system to investigate the role of chaos on the information gain in tomography madhok2014information ; sreeram2021quantum ; sahu2022effect ; PhysRevB.108.224306 , quantum simulations PhysRevLett.124.230501 and quantum phase transitions PhysRevA.106.032215 . Here, we connect two quantifiers of quantum chaos, namely Loschmidt echo and “error scrambling”, through continuous weak measurement tomography as illustrated in Fig. 1.
II Continuous measurement tomography
An ensemble of $N$ identical systems $\rho^{\otimes N}_{0}$ undergo a separable time evolution by a unitary $U(t)$. A weakly coupled probe will generate the measurement record by performing weak continuous measurement of an observable $\mathcal{O}$. For sufficiently weak coupling, the randomness of the measurement outcomes is dominated by the quantum noise in the probe rather than the measurement uncertainty, i.e., the projection noise. In this case, the quantum backaction is negligible, and the state remains approximately separable. Thus, we get the stochastic measurement record
@@ II Continuous measurement tomography
where ${\bf C=(\tilde{\mathcal{O}}^{T}\tilde{\mathcal{O}}})^{-1}$ is the covariance matrix and the inverse is Moore-Penrose pseudo inverse ben2003generalized in general. The estimated Bloch vector ${\bf r}_{ML}$ may not always represent a physical density matrix with nonnegative eigenvalues due to the finite signal-to-noise ratio. Therefore, we impose the constraint of positive semidefiniteness on the reconstructed density matrix and obtain the physical state closest to the maximum-likelihood estimate. To do this, we employ a convex optimization procedure vandenberghe1996semidefinite where the final estimate of the Bloch vector $\bf\bar{r}$ is obtained by minimizing the argument
@@
III Tomography with imperfect knowledge
The above description represents an ideal scenario where the experimentalist has complete knowledge of the true dynamics (which is symbolized as unprimed variables describing the observables, $\mathcal{O}_{n}$, and covariance matrix, $\bf C$, thus generated over time) and they can properly reconstruct the state using Eq. (4). However, in reality, one never knows the true underlying dynamics, and there is always a departure from the ideal case due to inevitable errors and perturbations to the true dynamics. Thus, the experimentalist, oblivious to these, models their estimation using a covariance matrix, ${\bf C^{\prime}=(\tilde{\mathcal{O^{\prime}}}^{T}\tilde{\mathcal{O^{\prime}}}})^{-1}$. Here, the primed variables represent the experimentalist’s knowledge of the dynamics in the laboratory, and as a result, they end up reconstructing an incorrect state as $\bar{\rho}^{\prime}$ from
@@ III Tomography with imperfect knowledge
Figure 2: Effect of perturbation on tomography quantified by different metrics as a function of time with an increase in the level of chaos. The kicked top Floquet map of true (perturbed) dynamics $U_{\tau}=e^{-i(\lambda+\delta\lambda)J^{2}_{z}/2J}e^{-i\alpha J_{x}}$ with $\delta\lambda=0.01$, and ideal (unperturbed) dynamics $U^{\prime}_{\tau}=e^{-i\lambda J^{2}_{z}/2J}e^{-i\alpha J_{x}}$ generate the time series of operators for a spin $j=10$ and fixed $\alpha=1.4$. (a) Average reconstruction fidelity $\langle\mathcal{F}\rangle$ of the state $\bar{\rho}^{\prime}$ derived from Eq. (6) relative to the actual state $\left|\psi_{0}\right\rangle$, where the average is taken over 100 Haar random states. (b) The operator Loschmidt echo $\mathcal{F}_{\mathcal{O}}$ between two operators. (c) The quantum relative entropy $\mathcal{D}_{KL}$ of the regularized operator evolved under unperturbed dynamics to the operator evolved under perturbed dynamics. (d) The operator incompatibility $\mathcal{I}_{\mathcal{O}}$ quantifies the scrambling of errors.
@@
Does having more chaos aid or inhibit quantum state reconstruction? As the chaos in the dynamics increases, the reconstruction fidelity of spin coherent states decreases sahu2022effect . This contrasts with results connecting information gain in tomography of random states with the degree of chaos in the dynamics that drives the system PhysRevLett.112.014102 ; madhok2016characterizing ; PhysRevA.104.032404 . The connection between chaos and information gain depends on the localization properties of the state, i.e. their inverse participation ratio, the degree of chaos, as well as how well the state is aligned with time-evolved measurement observables sahu2022effect . Therefore, to study the effect of the degree of chaos on the performance of noisy tomography purely, we consider random initial states measured via random initial observables (generated by rotating $J_{x}$ through random unitary) picked from the appropriate Haar measure. The connection between chaos and information gain depends on the localization properties of the state, i.e. their inverse participation ratio, the degree of chaos, as well as how well the state is aligned with time-evolved measurement observables sahu2022effect . Therefore, to study the effect of the degree of chaos on the performance of noisy tomography purely, we consider random initial states measured via random initial observables (generated by rotating $J_{x}$ through random unitary) picked from the appropriate Haar measure. We apply our reconstruction protocol on an ensemble of 100 random pure states sampled from the Haar measure on SU$(d)$, where $d=2j+1=21$. We choose one random initial observable and generate the measurement record from the repeated application of the Floquet map of the quantum kicked top. The fidelity of the reconstructed state $\bar{\rho}^{\prime}$ obtained from Eq. (6) is determined relative to the actual state $\left|\psi_{0}\right\rangle$, $\mathcal{F}=\left\langle\psi_{0}\right|\bar{\rho}^{\prime}\left|\psi_{0}\right\rangle$ as a function of time. We notice that the reconstruction fidelity increases in the beginning despite the errors, and after a certain period of time, it starts decaying. The rise in fidelity during the initial time period is because any information, even if partially inaccurate, about a completely unknown random state offsets the presence of errors in its estimation. However, as time progresses, the effect of errors becomes significant. Beyond a certain time, we observe a decline in fidelity as the dynamics continues to accumulate errors that dominate the archive of information present in the measurement record. Most interestingly, the rate of this fidelity decay is inversely correlated with the degree of chaos in the dynamics. This is the analog of an interplay between the rapid information gain due to Lyapunov divergence, a “quantum” analog of the classical KS entropy, and Loschmidt echo leading to errors that cause fidelity decay.
IV Quantum signatures of chaos in noisy tomography
@@
The operator Loschmidt echo that captures the overlap of the operators $\mathcal{O}_{n}$ and $\mathcal{O}^{\prime}_{n}$, decays with time. We can see from Fig. 2b that the operator Loschmidt echo decays much slower when the dynamics is chaotic than when it is regular. This behavior of the operator Loschmidt echo correlates positively with the rate of drop in reconstruction fidelity as demonstrated in Fig. 2a and Fig. 2b. The greater the distance between the operators at a given time, the greater the difference between the expectation values with respect to the state and the archive of the measurement record obtained through the time series. Our results give an operational interpretation of the operator Loschmidt echo by connecting it to a concrete physical task of continuous measurement quantum tomography. This also points to a beneficial way to probe these quantities in experiments using current techniques.
Another useful metric is the quantum relative entropy - a measure of distance between two quantum states. Here, we use this metric to measure the distance between two operators $\mathcal{O}_{n}$ and $\mathcal{O}^{\prime}_{n}$. To treat both observables as density operators, we regularize them as follows. We construct a positive operator from an observable by retaining its eigenvectors and taking the absolute value of its eigenvalues. To normalize this operator, we divide it by its trace. While calculating the relative entropy, we regularize the operator in the following way. First, we diagonalize the time-evolved operator $\mathcal{O}_{n}$ as
$\mathcal{O}_{n}=V_{n}D_{n}{V_{n}}^{\dagger},$ (8)
where $V_{n}$ is the unitary matrix that diagonalizes $\mathcal{O}_{n}$. In the second step, we take the absolute value of the eigenvalues of $D_{n}$ and divide it by its trace to get $\tilde{D}_{n}=|D_{n}|/{\rm{Tr}}\,(|D_{n}|)$. We, then, construct a positive operator with a unit trace that behaves as a density matrix while keeping the eigenvectors of the observable $\mathcal{O}_{n}$ intact as
$\rho_{\mathcal{O}_{n}}=V_{n}\tilde{D}_{n}{V_{n}}^{\dagger}.$ (9)
@@
This form of error scrambling is very general for dynamics with time-dependent as well as time-independent Hamiltonian. The operator incompatibility we have computed for the operators evolved through unperturbed and perturbed dynamics. This form is very general, and we do not assume any specific form of the error or any particular form of Hamiltonian (whether time-dependent or time-independent), as we can see below.
Let the operators at time $t$ be $\mathcal{O}(t)=U^{\dagger}(t)\mathcal{O}U(t)$ and $\mathcal{O^{\prime}}(t)=U^{\prime\dagger}(t)\mathcal{O}U^{\prime}(t)$, where $U(t)$ is the unitary for the unperturbed dynamics and $U^{\prime}(t)$ is the unitary for perturbed dynamics. Thus, the operator incompatibility is
@@
Now, one can simplify this expression using the Hermiticity property of the physical observables $\mathcal{O}(t)=\mathcal{O}^{\dagger}(t)$.
@@
We can further simplify the above expression by exploiting the cyclic property of trace operation and the property of a unitary operator $U^{\dagger}U=UU^{\dagger}=I$. Let us consider the first term in the above expression